# Gradient corrections to the local-equilibrium energy-momentum tensor in the Zubarev approach

### v2

@@ Abstract
We compute the expectation value of the energy-momentum tensor of a real scalar field in an approximation which accounts for space-time gradients of the hydrodynamical variables in local thermodynamical equilibrium. We show that the energy-momentum tensor receives corrections with respect to the standard local-equilibrium result. Notably, the relation between the energy density and pressure, i.e., the equation of state, is modified with respect to the one in global equilibrium. The obtained corrections might be relevant for systems created in relativistic hadron and heavy-ion collisions.
@@ I Introduction
For a system without a conserved charge which is in the hydrodynamic regime, instead of in global equilibrium, the relevant observable is the operator of the space-time dependent energy-momentum tensor $T^{\mu\nu}(x)$. A reduced description can be achieved (and hydrodynamic equations can be derived) utilizing Zubarev’s formalism of the non-equilibrium statistical operator Zubarev-1a ; Zubarev-1b ; Zubarev-1c ; Zubarev-1d (for modern developments, see Ref. Zubarev-2 and references therein). In Zubarev’s approach Zubarev-1a the non-equilibrium statistical operator maximizes the entropy subject to the following initial-state constraint imposed on a three-dimensional space-like hypersurface $\sigma(\tau_{0})$ with a time-like normal vector $n_{\mu}(x)$:
@@
The left-hand side of the above equation is determined by a true state of the system, with energy-momentum tensor $\tilde{T}^{\mu\nu}(x)$, and the right-hand side features the expectation value of the energy-momentum tensor operator calculated with the non-equilibrium statistical operator, $\langle\cdots\rangle\equiv{\rm Tr}[\rho_{\text{neq}}[\sigma(\tau_{0})]\cdots]$. Maximizing the entropy is essentially synonymous to assuming that the initial state of the system is in local thermodynamical equilibrium. Consequently, the non-equilibrium statistical operator reads Zubarev-3
@@
where $\mathrm{d}\sigma_{\mu}(y)\equiv\mathrm{d}\sigma\,n_{\mu}(y)$. The inverse four-temperature vector $\beta_{\nu}(y)\equiv\beta(y)u_{\nu}(y)$ is now a space-time dependent field, with $u_{\nu}(y)$ being the local four-velocity of the system (on a space-time point $y$ on $\sigma(\tau_{0})$), normalized as $u_{\nu}(y)u^{\nu}(y)=1$. In this approach there is a continuum of constraints: on the whole hypersurface $\sigma(\tau_{0})$ the relevant local observables, such as the energy-momentum tensor on the left-hand side of Eq. (7), are equal to the expectation values of the corresponding local quantum operators with respect to the statistical operator, i.e., the expectation value of the energy-momentum tensor operator on the right-hand side of Eq. (7).
The initial state, $\rho_{\text{neq}}[\sigma(\tau_{0})]$, characterizes an actual state of the system. This initial condition is not very restrictive if almost all initial states consistent with the above constraints evolve (in the Schrödinger picture) towards some hydrodynamical attractor (see, e.g., Ref. Spalinski ), where such states become empirically indistinguishable with respect to the set of relevant observables and most details of the actual initial microscopic conditions become irrelevant. In the original Zubarev approach the non-equilibrium statistical operator maximizes the entropy subject to constraints imposed in the infinitely remote past Zubarev-1a . For the description of the transient evolution of matter in hadron and heavy-ion collisions it is more natural, however, to use initial conditions which correspond to the beginning of the collision, i.e., on a suitably chosen hypersurface $\sigma(\tau_{0})$. The corresponding reformulation of Zubarev’s method was discussed in Ref. Zubarev-3 .
In the Heisenberg picture, by definition the statistical operator does not change with time or, if some one-parameter family of three-dimensional space-like hypersurfaces $\sigma(\tau)$ is defined, the statistical operator does not change with $\tau$. This allows to take space-time gradients out of the statistical averaging with $\rho_{\text{neq}}[\sigma(\tau_{0})]$, i.e., the conservation equations determining the Lagrange multipliers can be written as
@@
Note that Eq. (9) is time-reversible. [^3] In the original Zubarev approach Zubarev-1a , the reversibility of the exact conservation equations is broken by adding an infinitesimally small source term on the right-hand side of the evolution equation for the statistical operator in the Heisenberg representation.
Nevertheless, the fact that $\rho_{\text{neq}}[\sigma(\tau_{0})]$ does not change with time does not mean that it retains its functional form (8) on a three-dimensional space-like hypersurface $\sigma(\tau)$ with $\tau>\tau_{0}$. In fact, due to irreversible dissipative processes driven by space-time gradients of the hydrodynamical variables (i.e, in our case $\beta_{\nu}$), the system will deviate from the initial local-equilibrium state on $\sigma(\tau_{0})$ in the course of its evolution. This can be seen as follows. Using Gauss’ theorem and energy-momentum conservation one obtains
@@
Here, the four-dimensional space-time volume $\Omega$ is enclosed by the two space-like hypersurfaces $\sigma(\tau_{0})$ and $\sigma(\tau)$ and time-like hypersurfaces connecting these two, where $\beta_{\nu}(y)T^{\mu\nu}(y)$ is supposed to vanish. Then, the non-equilibrium statistical operator can be written as Zubarev-3
@@
and assume that $B$ is small compared to $A$, which is true if $\partial_{z,\mu}\beta_{\nu}(z)$ is sufficiently small. Consequently, to leading (zeroth) order in $B$ the non-equilibrium statistical operator is equal to the local-equilibrium statistical operator on the hypersurface $\sigma(\tau)$ Zubarev-3
@@
The transport coefficients are calculated in the Markovian (short-memory) approximation, which exploits the existence of disparate time-scales in the system. It is worth noting that this approximation is not only a useful tool for explicit calculations: The Markovian level of description mimics the effective loss of details about the initial conditions. Irreversible dissipative hydrodynamics appears as an effective theory of the slow degrees of freedom in its range of applicability, see, e.g., Refs. Zubarev-1a ; Zubarev-1b ; Zubarev-1c ; Zubarev-1d ; Zubarev-2 ; Zubarev-3 for derivations of dissipative hydrodynamical equations in the Zubarev approach.
@@
where $\epsilon(x)$ and $\mathcal{P}(x)$ are the local energy density and pressure of the fluid, respectively. However, this is not true, in fact $\langle T^{\mu\nu}(x)\rangle_{\text{leq}}\neq\tilde{T}^{\mu\nu}_{\text{id}}(x)$, due to the non-locality of $\rho_{\text{leq}}[\sigma(\tau)]$ introduced by the integration over the hypersurface $\sigma(\tau)$. In order to obtain the ideal-fluid form, one requires an additional approximation. Namely, one needs to assume that $\beta_{\mu}(y)$ varies sufficiently smoothly over the hypersurface $\sigma(\tau)$, such that the expectation value of some local operator $O(x)$, $\langle O(x)\rangle_{\text{leq}}={\rm Tr}[\rho_{\text{leq}}[\sigma(\tau)]O(x)]$, will be mainly determined by the value of the field $\beta_{\mu}(y)$ around the point $x$. Corrections can be systematically taken into account by expanding $\beta_{\mu}(y)$ in a Taylor series around $x$,
@@
If $C$ is a small correction to $\beta_{\nu}(x)P^{\nu}$, expectation values with respect to $\rho_{\text{leq}}[\sigma(\tau)]$ can be calculated perturbatively in a power series in $C$, just as the one in $B$ in Eq. (15) (see, e.g., Ref. Becattini-1 ). Since $C\sim T^{\mu\nu}$, to linear order the expectation value of the energy-momentum tensor receives corrections proportional to the two-point function of the energy-momentum tensor. In contrast to the terms $\sim B$ discussed previously, however, there are no traditional Kubo relations to relate these corrections to transport coefficients, since now there is no space-time integral over $\Omega$, but only an integral over the hypersurface $\sigma(\tau)$.
To leading (zeroth) order, we set $C=0$ and insert Eq. (19) into $\rho_{\text{leq}}[\sigma(\tau)]$. One then obtains an expression which is formally identical to the global-equilibrium statistical operator (6), but with a space-time dependent $\beta_{\nu}(x)$, which we refer to as “space-time ($x$-)dependent global equilibrium” ($x$-geq),
@@
For this reason, the space-time ($x$-)dependent global-equilibrium energy-momentum tensor is usually referred to as energy-momentum tensor in “local equilibrium”. We have seen that this is not quite correct, as the true local-equilibrium energy-momentum tensor, $\langle T^{\mu\nu}(x)\rangle_{\text{leq}}$, contains additional gradient terms $\sim C$ (and powers thereof). Therefore, although using “space-time dependent global equilibrium” sounds contradictory and much more cumbersome than “local equilibrium”, we will stick to this nomenclature in the following, since it accurately expresses the fact that expectation values are computed with the operator (20) instead of the local-equilibrium statistical operator (14).
@@
In this paper, we calculate the expectation value of the energy-momentum tensor operator of a real scalar field $\phi(x)$, using similar approximations for the non-equilibrium statistical operator as outlined above. However, we deviate from the above approach, which yields the well-known result (21), in one important aspect. Namely, the explicit expression for the energy-momentum tensor operator in general involves space-time gradients of the field operators. Using the stationarity of the non-equilibrium statistical operator (8), we can take these derivatives outside of the statistical averaging, similar as in Eq. (9). The approximations outlined above are then applied to the calculation of the statistical two-point function of the fields, $\langle\phi(x)\phi(y)\rangle$, rather than to the calculation of the expectation value $\langle T^{\mu\nu}(x)\rangle$ of the full energy-momentum tensor operator. We show that the expression obtained for $\langle T^{\mu\nu}(x)\rangle$ under these approximations involves corrections with respect to the ideal-fluid form (16). [^4] See also Refs. local ; local-2 . In particular, we demonstrate that the relation between pressure and energy density, i.e., the equation of state, is affected by these corrections and therefore is modified with respect to the case of global thermodynamical equilibrium. The magnitude of these corrections is determined by the ratio of the thermal wavelength $\lambda_{\text{th}}$ to the typical space-time homogeneity length $L$ of a given system. If the latter is much larger than the former, then these additional terms can be neglected. This is certainly true for the particular limit of global thermodynamical equilibrium, as well as for a coarse-grained description of macroscopic systems, where there is a clear separation between the microscopic and the macroscopic scales. However, this need not be the case, for example, for relativistic hadron and heavy-ion collisions, where small and highly inhomogeneous systems are created with a typical macroscopic time or length scale of a few femtometer.
@@
The remainder of this paper is structured as follows. In Sec. II we compute the expectation value of the energy-momentum tensor based on the idea outlined above, using a real scalar field as a simple, yet explicit example. In Sec. III we derive the corrections to the equation of state in global thermodynamical equilibrium arising from space-time gradients of the hydrodynamical fields. Our conclusions are given in Sec. IV.
@@ II Alternative calculation of the expectation value of the energy-momentum tensor
where the interacting part of the Lagrangian density, $\mathcal{L}_{\text{int}}(\phi)$, does not contain space-time derivatives. To calculate $\langle T^{\mu\nu}(x)\rangle$, one needs an explicit form for the operator of the energy-momentum tensor. The canonical one reads
@@
where the Lagrangian $\mathcal{L}$ is given by Eq. (22). Following the idea formulated in the introduction, one can take the space-time derivatives in Eq. (24) outside the statistical averaging, just as in Eq. (9), since the non-equilibrium statistical operator is stationary. We then obtain [^5] For simplicity, we will assume that the field expectation value vanishes, $\langle\phi(x)\rangle=0$.
@@
As advertised in the introduction, we now replace the statistical average with respect to the non-equilibrium statistical operator, $\langle\cdots\rangle=\mathrm{Tr}[\rho_{\text{neq}}[\sigma(\tau_{0})]\cdots]$, by the thermal average with respect to the space-time dependent global-equilibrium operator at point $z$, $\langle\cdots\rangle_{z-\text{geq}}=\mathrm{Tr}[\rho_{z-\text{geq}}(z)\cdots]$, i.e., we neglect all dissipative corrections $\sim B,C$. We will call the result the “gradient-corrected $z$-dependent global-equilibrium approximation” (gc-$z$-geq), which is then
@@
The terms in Eq. (32) with derivatives with respect to $z$ manifest deviations from (space-time dependent) global thermodynamical equilibrium where only relative positions matter. Therefore it is convenient to write the result as
@@
As we will now show, $\tilde{T}^{\mu\nu}_{z-\text{geq}}(z)$ coincides with the expectation value of the energy-momentum tensor calculated with the $z$-dependent global-equilibrium statistical operator (20), i.e., we will show that $\tilde{T}^{\mu\nu}_{z-\text{geq}}(z)=\tilde{T}^{\mu\nu}_{\text{id}}(z)$, in agreement with Eq. (21). To this end, let us Fourier-transform the statistical two-point function $F_{z-\text{geq}}\left(z-\frac{\Delta z}{2},z+\frac{\Delta z}{2}\right)$ with respect to the relative coordinate $\Delta z$,
@@
where $p_{0}=\sqrt{\mathbf{p}^{2}+m^{2}}$ is the on-shell energy of the particles and $\Delta_{\mu\nu}\equiv g_{\mu\nu}-u_{\mu}u_{\nu}$ is the projector onto the three-dimensional subspace orthogonal to $u^{\mu}$. Equations (38) and (39) are the standard expressions for the energy density and pressure of a non-interacting Bose gas of particles with mass $m$. We thus confirm that $\tilde{T}^{\mu\nu}_{z-\text{geq}}(z)=\tilde{T}^{\mu\nu}_{\text{id}}(z)$.
@@
III Pseudogauge dependence and gradient-corrected space-time dependent energy-momentum tensor
Equation (33) can be written in a more suggestive way as follows,

### v3

@@ Abstract
We compute the expectation value of the energy-momentum tensor of a real scalar field in an approximation which accounts for spacetime gradients of the hydrodynamical variables in local thermodynamical equilibrium. We show that the energy-momentum tensor receives corrections with respect to the standard local-equilibrium result. Notably, the relation between the energy density and pressure, i.e., the equation of state, is modified with respect to the one in global equilibrium. The obtained corrections might be relevant for systems created in relativistic hadron and heavy-ion collisions.
@@ I Introduction
For a system without a conserved charge which is in the hydrodynamic regime, instead of in global equilibrium, the relevant observable is the operator of the spacetime dependent energy-momentum tensor $T^{\mu\nu}(x)$. A reduced description can be achieved (and hydrodynamic equations can be derived) utilizing Zubarev’s formalism of the nonequilibrium statistical operator Zubarev-1a ; Zubarev-1b ; Zubarev-1c ; Zubarev-1d (for modern developments, see Ref. Zubarev-2 and references therein). In Zubarev’s approach Zubarev-1a the nonequilibrium statistical operator maximizes the entropy subject to the following initial-state constraint imposed on a three-dimensional spacelike hypersurface $\sigma(\tau_{0})$ with a timelike normal vector $n_{\mu}(x)$:
@@
The left-hand side of the above equation is determined by a true state of the system, with energy-momentum tensor $\tilde{T}^{\mu\nu}(x)$, and the right-hand side features the expectation value of the energy-momentum tensor operator calculated with the nonequilibrium statistical operator, $\langle\cdots\rangle\equiv{\rm Tr}[\rho_{\text{neq}}[\sigma(\tau_{0})]\cdots]$. Maximizing the entropy is essentially synonymous to assuming that the initial state of the system is in local thermodynamical equilibrium. Consequently, the nonequilibrium statistical operator reads Zubarev-3
@@
where $\mathrm{d}\sigma_{\mu}(y)\equiv\mathrm{d}\sigma\,n_{\mu}(y)$. The inverse four-temperature vector $\beta_{\nu}(y)\equiv\beta(y)u_{\nu}(y)$ is now a spacetime dependent field, with $u_{\nu}(y)$ being the local four-velocity of the system (on a spacetime point $y$ on $\sigma(\tau_{0})$), normalized as $u_{\nu}(y)u^{\nu}(y)=1$. In this approach there is a continuum of constraints: on the whole hypersurface $\sigma(\tau_{0})$ the relevant local observables, such as the energy-momentum tensor on the left-hand side of Eq. (7), are equal to the expectation values of the corresponding local quantum operators with respect to the statistical operator, i.e., the expectation value of the energy-momentum tensor operator on the right-hand side of Eq. (7).
The initial state, $\rho_{\text{neq}}[\sigma(\tau_{0})]$, characterizes an actual state of the system. This initial condition is not very restrictive if almost all initial states consistent with the above constraints evolve (in the Schrödinger picture) towards some hydrodynamical attractor (see, e.g., Ref. Spalinski ), where such states become empirically indistinguishable with respect to the set of relevant observables and most details of the actual initial microscopic conditions become irrelevant. In the original Zubarev approach the nonequilibrium statistical operator maximizes the entropy subject to constraints imposed in the infinitely remote past Zubarev-1a . For the description of the transient evolution of matter in hadron and heavy-ion collisions it is more natural, however, to use initial conditions which correspond to the beginning of the collision, i.e., on a suitably chosen hypersurface $\sigma(\tau_{0})$. The corresponding reformulation of Zubarev’s method was discussed in Ref. Zubarev-3 .
In the Heisenberg picture, by definition the statistical operator does not change with time or, if some one-parameter family of three-dimensional spacelike hypersurfaces $\sigma(\tau)$ is defined, the statistical operator does not change with $\tau$. This allows us to take space-time gradients out of the statistical averaging with $\rho_{\text{neq}}[\sigma(\tau_{0})]$, i.e., the conservation equations determining the Lagrange multipliers can be written as
@@
Note that Eq. (9) is time reversible. [^3] In the original Zubarev approach Zubarev-1a , the reversibility of the exact conservation equations is broken by adding an infinitesimally small source term on the right-hand side of the evolution equation for the statistical operator in the Heisenberg representation.
Nevertheless, the fact that $\rho_{\text{neq}}[\sigma(\tau_{0})]$ does not change with time does not mean that it retains its functional form (8) on a three-dimensional spacelike hypersurface $\sigma(\tau)$ with $\tau>\tau_{0}$. In fact, due to irreversible dissipative processes driven by spacetime gradients of the hydrodynamical variables (i.e, in our case $\beta_{\nu}$), the system will deviate from the initial local-equilibrium state on $\sigma(\tau_{0})$ in the course of its evolution. This can be seen as follows: using Gauss’ theorem and energy-momentum conservation one obtains
@@
Here, the four-dimensional spacetime volume $\Omega$ is enclosed by the two spacelike hypersurfaces $\sigma(\tau_{0})$ and $\sigma(\tau)$ and timelike hypersurfaces connecting these two, where $\beta_{\nu}(y)T^{\mu\nu}(y)$ is supposed to vanish. Then, the nonequilibrium statistical operator can be written as Zubarev-3
@@
and assume that $B$ is small compared to $A$, which is true if $\partial_{z,\mu}\beta_{\nu}(z)$ is sufficiently small. Consequently, to leading (zeroth) order in $B$ the nonequilibrium statistical operator is equal to the local-equilibrium statistical operator on the hypersurface $\sigma(\tau)$ Zubarev-3
@@
The transport coefficients are calculated in the Markovian (short-memory) approximation, which exploits the existence of disparate time scales in the system. It is worth noting that this approximation is not only a useful tool for explicit calculations: the Markovian level of description mimics the effective loss of details about the initial conditions. Irreversible dissipative hydrodynamics appears as an effective theory of the slow degrees of freedom in its range of applicability, see, e.g., Refs. Zubarev-1a ; Zubarev-1b ; Zubarev-1c ; Zubarev-1d ; Zubarev-2 ; Zubarev-3 for derivations of dissipative hydrodynamical equations in the Zubarev approach.
@@
where $\epsilon(x)$ and $\mathcal{P}(x)$ are the local energy density and pressure of the fluid, respectively. However, this is not true, in fact $\langle T^{\mu\nu}(x)\rangle_{\text{leq}}\neq\tilde{T}^{\mu\nu}_{\text{id}}(x)$, due to the nonlocality of $\rho_{\text{leq}}[\sigma(\tau)]$ introduced by the integration over the hypersurface $\sigma(\tau)$. In order to obtain the ideal-fluid form, one requires an additional approximation. Namely, one needs to assume that $\beta_{\mu}(y)$ varies sufficiently smoothly over the hypersurface $\sigma(\tau)$, such that the expectation value of some local operator $O(x)$, $\langle O(x)\rangle_{\text{leq}}={\rm Tr}[\rho_{\text{leq}}[\sigma(\tau)]O(x)]$, will be mainly determined by the value of the field $\beta_{\mu}(y)$ around the point $x$. Corrections can be systematically taken into account by expanding $\beta_{\mu}(y)$ in a Taylor series around $x$,
@@
If $C$ is a small correction to $\beta_{\nu}(x)P^{\nu}$, expectation values with respect to $\rho_{\text{leq}}[\sigma(\tau)]$ can be calculated perturbatively in a power series in $C$, just as the one in $B$ in Eq. (15) (see, e.g., Ref. Becattini-1 ). Since $C\sim T^{\mu\nu}$, to linear order the expectation value of the energy-momentum tensor receives corrections proportional to the two-point function of the energy-momentum tensor. In contrast to the terms $\sim B$ discussed previously, however, there are no traditional Kubo relations to relate these corrections to transport coefficients, since now there is no spacetime integral over $\Omega$, but only an integral over the hypersurface $\sigma(\tau)$.
To leading (zeroth) order, we set $C=0$ and insert Eq. (19) into $\rho_{\text{leq}}[\sigma(\tau)]$. One then obtains an expression which is formally identical to the global-equilibrium statistical operator (6), but with a spacetime dependent $\beta_{\nu}(x)$, which we refer to as “spacetime ($x$-)dependent global equilibrium” ($x$-geq),
@@
For this reason, the spacetime ($x$-)dependent global-equilibrium energy-momentum tensor is usually referred to as energy-momentum tensor in “local equilibrium”. We have seen that this is not quite correct, as the true local-equilibrium energy-momentum tensor, $\langle T^{\mu\nu}(x)\rangle_{\text{leq}}$, contains additional gradient terms $\sim C$ (and powers thereof). Therefore, although using “spacetime dependent global equilibrium” sounds contradictory and much more cumbersome than “local equilibrium”, we will stick to this nomenclature in the following, since it accurately expresses the fact that expectation values are computed with the operator (20) instead of the local-equilibrium statistical operator (14).
@@
In this paper, we calculate the expectation value of the energy-momentum tensor operator of a real scalar field $\phi(x)$, using similar approximations for the nonequilibrium statistical operator as outlined above. However, we deviate from the above approach, which yields the well-known result (21), in one important aspect. Namely, the explicit expression for the energy-momentum tensor operator in general involves spacetime gradients of the field operators. Using the stationarity of the nonequilibrium statistical operator (8), we can take these derivatives outside of the statistical averaging, as in Eq. (9). The approximations outlined above are then applied to the calculation of the statistical two-point function of the fields, $\langle\phi(x)\phi(y)\rangle$, rather than to the calculation of the expectation value $\langle T^{\mu\nu}(x)\rangle$ of the full energy-momentum tensor operator. We show that the expression obtained for $\langle T^{\mu\nu}(x)\rangle$ under these approximations involves corrections with respect to the ideal-fluid form (16). [^4] See also Refs. local ; local-2 . In particular, we demonstrate that the relation between pressure and energy density, i.e., the equation of state, is affected by these corrections and therefore is modified with respect to the case of global thermodynamical equilibrium. The magnitude of these corrections is determined by the ratio of the thermal wavelength $\lambda_{\text{th}}$ to the typical spacetime homogeneity length $L$ of a given system. If the latter is much larger than the former, then these additional terms can be neglected. This is certainly true for the particular limit of global thermodynamical equilibrium, as well as for a coarse-grained description of macroscopic systems, where there is a clear separation between the microscopic and the macroscopic scales. However, this need not be the case, for example, for relativistic hadron and heavy-ion collisions, where small and highly inhomogeneous systems are created with a typical macroscopic time or length scale of a few femtometers.
@@
The remainder of this paper is structured as follows: in Sec. II we compute the expectation value of the energy-momentum tensor based on the idea outlined above, using a real scalar field as a simple, yet explicit example. In Sec. III we derive the corrections to the equation of state in global thermodynamical equilibrium arising from spacetime gradients of the hydrodynamical fields. Our conclusions are given in Sec. IV.
@@ II Alternative calculation of the expectation value of the energy-momentum tensor
where the interacting part of the Lagrangian density, $\mathcal{L}_{\text{int}}(\phi)$, does not contain spacetime derivatives. To calculate $\langle T^{\mu\nu}(x)\rangle$, one needs an explicit form for the operator of the energy-momentum tensor. The canonical one reads
@@
where the Lagrangian $\mathcal{L}$ is given by Eq. (22). Following the idea formulated in the introduction, one can take the spacetime derivatives in Eq. (24) outside the statistical averaging, just as in Eq. (9), since the nonequilibrium statistical operator is stationary. We then obtain [^5] For simplicity, we will assume that the field expectation value vanishes, $\langle\phi(x)\rangle=0$.
@@
As advertised in the introduction, we now replace the statistical average with respect to the nonequilibrium statistical operator, $\langle\cdots\rangle=\mathrm{Tr}[\rho_{\text{neq}}[\sigma(\tau_{0})]\cdots]$, by the thermal average with respect to the spacetime dependent global-equilibrium operator at point $z$, $\langle\cdots\rangle_{z-\text{geq}}=\mathrm{Tr}[\rho_{z-\text{geq}}(z)\cdots]$, i.e., we neglect all dissipative corrections $\sim B,C$. We will call the result the “gradient-corrected $z$-dependent global-equilibrium approximation” (gc-$z$-geq), which is then
@@
The terms in Eq. (32) with derivatives with respect to $z$ manifest deviations from (spacetime dependent) global thermodynamical equilibrium where only relative positions matter. Therefore it is convenient to write the result as
@@
As we will now show, $\tilde{T}^{\mu\nu}_{z-\text{geq}}(z)$ coincides with the expectation value of the energy-momentum tensor calculated with the $z$-dependent global-equilibrium statistical operator (20), i.e., we will show that $\tilde{T}^{\mu\nu}_{z-\text{geq}}(z)=\tilde{T}^{\mu\nu}_{\text{id}}(z)$, in agreement with Eq. (21). To this end, let us Fourier-transform the statistical two-point function $F_{z-\text{geq}}\left(z-\frac{\Delta z}{2},z+\frac{\Delta z}{2}\right)$ with respect to the relative coordinate $\Delta z$, as follows:
@@
where $p_{0}=\sqrt{\mathbf{p}^{2}+m^{2}}$ is the on shell energy of the particles and $\Delta_{\mu\nu}\equiv g_{\mu\nu}-u_{\mu}u_{\nu}$ is the projector onto the three-dimensional subspace orthogonal to $u^{\mu}$. Equations (38) and (39) are the standard expressions for the energy density and pressure of a noninteracting Bose gas of particles with mass $m$. We thus confirm that $\tilde{T}^{\mu\nu}_{z-\text{geq}}(z)=\tilde{T}^{\mu\nu}_{\text{id}}(z)$.
@@
III Pseudogauge dependence and gradient-corrected spacetime dependent energy-momentum tensor
Equation (33) can be written in a more suggestive way as follows: